# UniHR: Hierarchical Representation Learning for Unified Knowledge Graph Link Prediction

### v6

@@ Abstract
Real-world knowledge graphs (KGs) contain not only standard triple-based facts, but also more complex, heterogeneous types of facts, such as hyper-relational facts with auxiliary key-value pairs, temporal facts with additional timestamps, and nested facts that imply relationships between facts. These richer forms of representation have attracted significant attention due to their enhanced expressiveness and capacity to model complex semantics in real-world scenarios. However, most existing studies suffer from two main limitations: (1) they typically focus on modeling only specific types of facts, thus making it difficult to generalize to real-world scenarios with multiple fact types; and (2) they struggle to achieve generalizable hierarchical (inter-fact and intra-fact) modeling due to the complexity of these representations. To overcome these limitations, we propose UniHR, a Unified Hierarchical Representation learning framework, which consists of a learning-optimized Hierarchical Data Representation (HiDR) module and a unified Hierarchical Structure Learning (HiSL) module. The HiDR module unifies hyper-relational KGs, temporal KGs, and nested factual KGs into triple-based representations. Then HiSL incorporates intra-fact and inter-fact message passing, focusing on enhancing both semantic information within individual facts and enriching the structural information between facts. To go beyond the unified method itself, we further explore the potential of unified representation in complex real-world scenarios. Extensive experiments on 9 datasets across 5 types of KGs demonstrate the effectiveness of UniHR and highlight the strong potential of unified representations. Code and data are available at https://github.com/zjukg/UniHR.
@@ Introduction
To fill this research gap, we propose UniHR, a Unified Hierarchical Representation learning framework, which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module. HiDR module standardizes hyper-relational facts, nested factual facts, and temporal facts into the form of triples without loss of information. Furthermore, HiSL module captures local semantic information during intra-fact message passing and then utilizes inter-fact message passing to enrich the global structure information to obtain better node embeddings based on HiDR form. Finally, the updated embeddings are fed into decoders for link prediction. Apart from the unification of method itself, UniHR’s unified representation enables flexible extensions. Unlike previous KG-specific models, UniHR accommodates more complex scenarios, such as compositional knowledge graphs, multi-task learning, and joint training on hybrid facts, thereby paving the way for pre-trained models across diverse KG types. Empirical results demonstrate the UniHR achieves competitive performance on most metrics without excessive increase in computational cost and highlight the strong potential of unified representations: Our contributions can be summarized as follows.
@@
This section discusses the limitations of existing methods in hierarchical modeling.
@@
For numerical atomic nodes, namely timestamps in temporal knowledge graphs, we encode the timestamp $\tau$ into an embedding with Time2Vec (Time2Vec):
$\mathbf{h}_{\tau}=\omega_{p}\sin\left(f_{p}(\tau)\right)+f_{np}(\tau),$ (2)
where $f_{p}\text{:}\;\mathbb{R}^{1}\rightarrow\mathbb{R}^{d}$ and $f_{np}\text{:}\;\mathbb{R}^{1}\rightarrow\mathbb{R}^{d}$ are both 1-layer MLPs as periodic and non-periodic functions, and $\omega_{p}\in\mathbb{R}^{1}$ is a learnable parameter for scaling the periodic features.
@@
For HKG, we compare with NaLP (NaLP), StarE (stare), GRAN (gran), tNaLP (tnalp), HyNT (hynt), ShrinkE (shrinke), HAHE (hahe), NYLON (nylon) and HyperSAT (HyperSAT). For NKG, QuatE (quate), BiQUE (bique), BiVE (bive), NestE (neste), HOKE (HOKE) and GRADATE (GRADATE) are chosen as baselines. For TKG, we compare against following methods: ATiSE (ATiSE), TGeomE+ (TGeomE+), HGE (HGE), HyperSAT (HyperSAT), DuaTHP (duathp), ECEformer (eceformer) and 5EL (5EL).
@@
Variant FBHE (NKG) DB15K (HKG) wikidata12k (TKG)
@@
To analyze the contribution of different modules across various KG types, we present ablation results in Table 4. It can be observed that both intra-fact and inter-fact message passing contribute to performance improvement. In particular, intra-fact message passing proves to be more beneficial for NKGs. We attribute this to fact nodes in NKGs being inherently composed of other atomic nodes, making triple prediction rely heavily on comprehensive bi-level fact semantics. In contrast, HKGs and TKGs focus solely on atomic nodes, whose representations are not dependent on other nodes. Therefore, inter-fact message passing, by capturing the global context among facts, works more effective for HKGs and TKGs, leading to further performance improvements.
@@ Conclusion
In this paper, we propose UniHR, a unified hierarchical KG representation learning framework consisting of a learning-optimized Hierarchical Data Representation (HiDR) module and a unified Hierarchical Structure Learning (HiSL) module. The HiDR module unifies hyper-relational, nested and temporal facts into the triple form. Moreover, HiSL captures local semantic information within facts and global structural information between facts. Extensive experiments show UniHR achieves the best or competitive performance across 5 types of KGs over 9 datasets and further highlight the strong potential of unified representations across 3 complex scenarios. We hope UniHR will encourage research on unified KG representation learning and facilitate the development of pre-trained models for real-world KGs with heterogeneous fact types.

### v7

@@ Abstract
Real-world knowledge graphs (KGs) contain not only standard triple-based facts, but also more complex, heterogeneous types of facts, such as hyper-relational facts with auxiliary key-value pairs, temporal facts with additional timestamps, and nested facts that imply relationships between facts. These richer forms of representation have attracted significant attention due to their enhanced expressiveness and capacity to model complex semantics in real-world scenarios. However, most existing studies suffer from two main limitations: (1) they typically focus on modeling only specific types of facts, thus making it difficult to generalize to real-world scenarios with multiple fact types; and (2) they struggle to achieve generalizable hierarchical (inter-fact and intra-fact) modeling due to the complexity of these representations. To overcome these limitations, we propose UniHR, a Unified Hierarchical Representation learning framework, which consists of a learning-optimized Hierarchical Data Representation (HiDR) module and a unified Hierarchical Structure Learning (HiSL) module. The HiDR module unifies hyper-relational KGs, temporal KGs, and nested factual KGs into triple-based representations. Then HiSL incorporates intra-fact and inter-fact message passing, focusing on enhancing both semantic information within individual facts and enriching the structural information between facts. To go beyond the unified method itself, we further explore the potential of unified representation in complex real-world scenarios. Extensive experiments on 9 datasets across 5 types of KGs demonstrate the effectiveness of UniHR and highlight the strong potential of unified representations.
Codehttps://github.com/zjukg/UniHR
@@ Introduction
To fill this research gap, we propose UniHR, a Unified Hierarchical Representation learning framework, which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module. HiDR module standardizes hyper-relational facts, nested factual facts, and temporal facts into the form of triples without loss of information. Furthermore, HiSL module captures local semantic information during intra-fact message passing and then utilizes inter-fact message passing to enrich the global structure information to obtain better node embeddings based on HiDR form. Finally, the updated embeddings are fed into decoders for link prediction. Apart from the unification of method itself, UniHR’s unified representation enables flexible extensions. Unlike previous KG-specific models, UniHR accommodates more complex scenarios, such as compositional knowledge graphs, multi-task learning, and joint training on hybrid facts, thereby paving the way for pre-trained models across diverse KG types. Our contributions can be summarized as follows.
@@
For numerical atomic nodes, namely timestamps in temporal knowledge graphs, we encode the timestamp $\tau$ into an embedding with Time2Vec (Time2Vec): $\mathbf{h}_{\tau}=\omega_{p}\sin\left(f_{p}(\tau)\right)+f_{np}(\tau)$, where $f_{p}\text{:}\;\mathbb{R}^{1}\rightarrow\mathbb{R}^{d}$ and $f_{np}\text{:}$$\mathbb{R}^{1}\rightarrow\mathbb{R}^{d}$ are both 1-layer linear layers as periodic and non-periodic functions, and $\omega_{p}\in\mathbb{R}^{1}$ is a learnable parameter for scaling the periodic features.
@@
For HKG, we compare with NaLP (NaLP), StarE (stare), GRAN (gran), tNaLP (tnalp), HyNT (hynt), ShrinkE (shrinke), HAHE (hahe), NYLON (nylon) and HyperSAT (HyperSAT). For NKG, QuatE (quate), BiQUE (bique), BiVE (bive), NestE (neste), HOKE (HOKE) and GRADATE (GRADATE) are chosen as baselines. For TKG, we compare against following methods: ATiSE (ATiSE), TGeomE+ (TGeomE+), HGE (HGE), DuaTHP (duathp), ECEformer (eceformer) and 5EL (5EL).
@@
Variant FBHE (N) DB15K (H) wikidata12k (T)
@@
To analyze the contribution of different modules across various KG types, we present ablation results in Table 4. It can be observed that both intra-fact and inter-fact message passing contribute to performance improvement. In particular, intra-fact message passing proves to be more beneficial for NKGs. We attribute this to fact nodes in NKGs being inherently composed of other atomic nodes, making triple prediction rely heavily on comprehensive bi-level fact semantics. In contrast, HKGs and TKGs focus solely on atomic nodes, whose representations are not dependent on other nodes. Therefore, inter-fact message passing, by capturing the global context among facts, works more effective for HKGs and TKGs, leading to better performances.
@@ Conclusion
In this paper, we propose UniHR, a unified hierarchical KG representation learning framework consisting of a learning-optimized Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module. The HiDR module unifies hyper-relational, nested and temporal facts into the triple form. Moreover, HiSL captures local semantic information within facts and global structural information between facts. Extensive experiments show UniHR achieves the best or competitive performance across 5 types of KGs over 9 datasets and further highlight the strong potential of unified representations across 3 complex scenarios.